OpenStudy (dls):

Integrate cos^6 x using complex numbers?

OpenStudy (dls):

like.. z=cosx+isinx \[\frac{1}{z}=\cos x-i \sin x\] \[(z+\frac{1}{z})^6=(2cosx)^6\] now we use binomial..dont know how to do this part..

OpenStudy (dls):

\[\LARGE C(6,0) (z^6+\frac{1}{z^6})+C(6,1)..\] i basically want to know what to put in place of z^6+1/z^6 etc

OpenStudy (dumbcow):

\[\cos^{6} x = \frac{1}{2^{6}}(z+z^{-1})^{6}\] \[(z+z^{-1})^{6} = z^{6}+6z^{4}+15z^{2}+20+15z^{-2}+6z^{-4}+z^{-6}\] also \[z = e^{ix}\]

hartnn (hartnn):

z^6 = cos 6x + i sin6x 1/z^6 = cos 6x - i sin 6x z^6 +1/z^6 = 2 cos 6x

OpenStudy (dls):

what now?

hartnn (hartnn):

z^4+1/z^4 = 2 cos 4x

hartnn (hartnn):

z^2+1/z^2 = 2 cos 2x

hartnn (hartnn):

you have the form A cos 6x +Bcos4x+Ccos 2x so easy to integrate

OpenStudy (dls):

oh! thanks!

hartnn (hartnn):

+20 ofcourse welcome ^_^

OpenStudy (anonymous):

@hartnn +2cosx i sinx ?

OpenStudy (anonymous):

in z^6 = cos 6x + i sin6x